# Softplus Attention with Re-weighting Boosts Length Extrapolation in Large Language Models

### v2

@@ Abstract
Large language models have achieved remarkable success in recent years, primarily due to the implementation of self-attention mechanisms. However, traditional Softmax attention suffers from numerical instability and reduced performance as the length of inference tokens increases. This paper addresses these issues by decomposing the Softmax operation into a non-linear transformation and the $l_{1}$-norm. We identify the latter as essential for maintaining model performance. By replacing the non-linear transformation with the Softplus activation function and introducing a dynamic scale factor for different token lengths based on invariance entropy, we create a novel attention mechanism with performance better than conventional Softmax attention across various inference lengths. To further improve the length extrapolation ability of the proposed attention mechanism, we introduce a fine-tuning-free re-weighting mechanism that amplifies significant attention weights while diminishing weaker ones, enabling the model to concentrate more effectively on relevant tokens without requiring retraining. When combined with our proposed attention mechanism, this approach demonstrates significant promise in managing longer sequences, maintaining nearly constant validation loss even at 16$\times$ the training token length while ensuring numerical stability. Our code is available at: https://github.com/iminfine/freeatten.
Large Language Models; Softmax-free attention; Length Extrapolation
1 Introduction
Large Language Models (LLMs) have achieved significant success in recent years, largely due to the use of self-attention mechanism. This approach enables models to assess the importance of different words within a sentence, capturing complex relationships and dependencies in the data. Its effectiveness is mainly attributed to the Softmax operation. When the Softmax operation is omitted or replaced with alternative activation functions, model performance tends to decline (Wortsman et al., 2023; Ramapuram et al., 2024; Shen et al., 2023). Although widely used, the Softmax self-attention has two main limitations. Firstly, it suffers from numerical instability due to the natural exponential function, especially when scaling model sizes to trillions of parameters (sta, ). Secondly, as the token length increases during inference, the attention scores calculated by self-attention become smoother and lack distinct peaks. This hinders the model’s ability to establish connections between relevant tokens, thereby affecting the length extrapolation capabilities of transformers (Chiang & Cholak, 2022). This paper addresses these two problems simultaneously: developing a Softmax-free attention mechanism that offers improved numerical stability and better performance even at large token lengths.
@@ 1 Introduction
Softmax calculates row attention scores in two steps: first, the scores are converted to a positive domain using the natural exponential function, and then each row is normalised to ensure the scores sum to one. This process inherently produces non-negative outputs, a trait long deemed vital for its success in attention mechanisms (Tsai et al., 2019; Katharopoulos et al., 2020; Han et al., 2023; Arora et al., 2024). However, this assumption has come under scrutiny, as studies substituting Softmax with other non-negative functions (Ramapuram et al., 2024; Li et al., 2022; Shen et al., 2023; Wortsman et al., 2023) report performance declines. Our experiments further explore this by assessing the necessity of non-negative attention scores, revealing that while non-negativity isn’t crucial, the normalisation process is key to sustaining performance.
Inspired by these observations, we decompose the Softmax operation into a non-linear transformation (the natural exponential function) and the $l_{1}$-norm, concluding that the $l_{1}$-norm is crucial to the success of traditional attention mechanisms. We then replace the non-linear transformation with various activation functions, resulting in models that match the performance of the original Softmax-based model. Furthermore, by substituting the non-linear transformation with the Softplus activation function and introducing a dynamic length scale factor that accounts for different token lengths based on invariance entropy, we propose a novel Length Scaled Softplus Attention (LSSA) mechanism. Our experiments demonstrate that LSSA outperforms the standard attention mechanism not only at the training sequence length but also for longer sequences. This highlights the potential of the proposed method to enhance the performance of LLMs.
To ensure that self-attention produces distinct peaks even for very long token length, we introduce a novel fine-tuning-free re-weighting approach that enhances the model’s capacity to concentrate on relevant tokens while filtering out noise. By applying a power transformation to the attention scores and $l_{1}$-norm afterwards, this re-weighting mechanism effectively amplifies stronger attention weights and suppresses weaker ones without retraining or architectural changes. Empirical results demonstrate that this re-weighting mechanism can be seamlessly integrated with various attention variants, leading to improved performance. When combined with our proposed LSSA, as shown in Figure 1 it shows particular promise in managing longer sequences, maintaining nearly constant validation loss even at 16 $\times$ the training token length while ensuring numerical stability.
@@
  1. 1.
    We decompose the Softmax operation into a non-linear transformation and the $l_{1}$-norm and find the normalisation process is key to sustaining performance. By replacing the non-linear transformation with various activation functions, we develop multiple attention variants that are competitive with Softmax.
  2. 2.
    We introduce a novel Softmax-free attention mechanism, referred to as LSSA, which incorporates a dynamic length scale factor based on invariance entropy, demonstrating superior performance compared to standard attention mechanisms across a variety of sequence lengths.
  3. 3.
    We propose a novel fine-tuning-free re-weighting approach that enhances the model’s ability to focus on relevant tokens by amplifying significant attention weights while reducing weaker ones without requiring retraining or architectural modifications. When combined with our proposed LSSA, this re-weighted model (LSSAR) excels particularly in length extrapolation while ensuring numerical stability.
2 Method
2.1 Softmax Decomposition
Scaled dot-product attention transforms queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) through a two-step procedure. First, attention scores $\mathbf{A}$ are produced via:
@@ 2.1 Softmax Decomposition
Here, $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times d}$, where $L$ represents the sequence length and $d$ denotes the dimensionality. The optional term $\mathbf{M}$ acts as a mask with shape $L\times L$ , which is particularly beneficial in causal self-attention contexts. The mask $\mathbf{M}$ is constructed with zeros below the diagonal and $\infty$- ∞ above it, ensuring that the attention mechanism only considers past and present tokens, thereby preserving the autoregressive nature of the model. The attention scores are subsequently utilised to compute the outputs as $\mathbf{O}=\mathbf{A}\mathbf{V}$.
Operation Softmax Inverse Re-centered Re-centered & $l_{1}$-norm
Loss 3.1911 3.1915 3.2008 3.1954
The Softmax function is a crucial component of the attention mechanism. It transforms a vector of raw scores $\mathbf{x}$ into a probability distribution, where the probabilities of each token sum to one. The outputs of Softmax are inherently non-negative, a feature considered essential for the success of Softmax attention. However, this assumption is somewhat questionable, as empirical studies that replaced the Softmax function with other non-negative output activation functions have shown a decline in model performance.
To investigate this further, we conducted experiments to test the necessity of non-negative attention scores. As illustrated in Table 1, one approach involved inverting the Softmax outputs by multiplying them by negative one, converting them to non-positive values. This inversion resulted in the validation loss remaining nearly unchanged, suggesting that non-negativity is not crucial for maintaining LLM performance. Additionally, we performed two more experiments. The first involved re-centring the attention scores. For each row of attention scores, we subtracted the mean value of that row, ensuring that each row had a mean of zero and, thus, contained both negative and positive values [^1] To prevent training instability due to the absence of positive values in the initial rows, the first three rows were left unmodified.. Inspired by the normalisation approach used in the Softmax operation, the second experiment further applied $l_{1}$-normalisation to each row, ensuring that the sum of the absolute values equaled one.
The results indicate that non-negative attention scores are not necessary for model performance. However, the $l_{1}$-norm is beneficial for maintaining performance. Motivated by these observations, we decomposed the Softmax operation into a non-linear transformation, the exponential function, followed by the $l_{1}$-norm:
$\operatorname{Softmax}(\mathbf{x})$ $=\frac{e^{\mathbf{x}}}{\sum_{j}e^{x_{j}}}=\frac{e^{\mathbf{x}}}{\|e^{\mathbf{x}}\|_{1}}=\frac{\phi(\mathbf{x})}{\|\phi(\mathbf{x})\|_{1}},$ (2)
[…]
Thus, the original attention function (Equation 1) can be expressed in a more general form:
@@
The subscripts $i$ and $j$ represent the row and column indices respectively. $\otimes$ denotes element-wise multiplication. Different from $\mathbf{M}$ defined in Equation 1, the mask $\mathbf{M^{\prime}}$ here is constructed with ones below the diagonal and zeros above it. It is evident that the original attention function (Equation 1) is a special case of Equation 3: one where the non-linear transformation function $\phi$ is the exponential function. This decomposition retains the essential characteristics of the attention mechanism while allowing for greater flexibility in the choice of activation functions. By replacing $\phi$ with various activation functions, we develop alternative attention mechanisms.
2.2 Length Scaled Softplus Attention
@@
In this subsection, we present Length Scaled Softplus Attention (LSSA), a novel attention variant designed to enhance the scalability and performance of LLMs. Building upon the foundational work in attention mechanisms, LSSA introduces a novel scaling factor that accounts for both the sequence length and the model’s dimensionality, thereby addressing limitations associated with traditional attention methods in handling long sequences during inference.
@@
LSSA utilises the Softplus activation function to introduce non-linearity while maintaining smooth gradients essential for stable training dynamics. Through empirical testing, we evaluated several widely-adopted activation functions by substituting them for $\phi$ into Equation 3. Our experiments in Section 3.2 revealed that Softplus outperforms the traditional exponential function ($e^{x}$), thereby justifying its adoption in the design of LSSA.
@@
In contrast to scaled dot-product attention, cosine similarity attention employs the $l$2-norm for each row of $\mathbf{Q}$ and $\mathbf{K}$. This approach has been shown to produce more moderate attention weights, which can enhance performance across various tasks (Henry et al., 2020; Dehghani et al., 2023; Liu et al., 2022). However, in high-dimensional spaces, as the number of dimensions increases, two randomly selected vectors are likely to become orthogonal. This phenomenon causes the elements of the product of normalised $\|\mathbf{Q}\|_{2}$ and $\|\mathbf{K}\|_{2}^{T}$ to approach zero, thereby worsening the vanishing gradient problem. Consequently, a scale factor becomes essential for cosine similarity attention, and this factor should be associated with the dimensionality $d$.
@@
Previous work demonstrated that replacing the traditional scaling factor $1/\sqrt{d}$ with $\log{L}/\sqrt{d}$ in scaled dot-product attention enhances the length extrapolation capabilities of transformers (Chiang & Cholak, 2022). Furthermore, (Su, 2021) highlighted that the inclusion of the $\log{L}$ factor aids in maintaining entropy invariance with different token length, thereby facilitating better extrapolation to unknown sequence lengths. Extending these findings, our approach adjusts the scaling factor to a dynamic length scale factor, reflecting the variation in token length across different rows of attention scores. Specifically, we set the scaling factor to $\log{d}\log{\mathbf{N}}$, where $\mathbf{N}$ is an $L\times L$ matrix, with each row containing identical elements corresponding to the number of tokens involved in attention for the $i$-th row, i.e., the number of tokens without masking. Clearly, $\mathbf{N}_{ij}=i$. This ensures that the attention mechanism remains robust across varying sequence lengths and model dimensions.
The formulation of LSSA is mathematically defined as:
@@
2.3 Attention Re-weighting Mechanism
While scaled dot-product attention has achieved remarkable success in LLMs, it inherently activates all input tokens to compute attention scores. This approach is somewhat counter-intuitive, as tokens positioned at the beginning of the input sequence are less likely to be relevant to tokens along the diagonal of row attention scores. Consequently, early tokens should ideally have minimal or no influence as $i$ increases when generating attention scores.
Experiments with ReLU-based attention, discussed in Section 3.2, demonstrate a minor performance decrease compared to the original attention mechanism. The observed performance decrement can be attributed to the distribution of positive values within the row attention scores. Specifically, positive values are predominantly concentrated around the diagonal, while other positions remain negative. The ReLU activation function tends to suppress these negative inputs, resulting in a limited number of activated neurons during training. This sparsity restricts the model’s capacity to capture essential relationships across the entire sequence, thereby adversely impacting overall performance. These results support the need for a re-weighting mechanism.
This phenomenon cannot be adequately addressed by traditional normalisation methods such as LayerNorm (Ba et al., 2016), as LayerNorm does not alter the underlying distribution of row attention scores. One potential solution is to transform all attention score values to the positive domain using functions like $e^{x}$ or Softplus before applying LayerNorm. As all values become positive through such transformations, employing the $l_{1}$-norm emerges as a more effective strategy. The $l_{1}$-norm is advantageous because it is simple to calculate and ensures that the mean value of the effective values of each row (the positive values that sum to one) remains constant (equals to $\frac{1}{i}$), thereby maintaining the integrity of the attention distribution. In this context, the Softmax operation in classic attention mechanism and Softplus with $l_{1}$-norm in LSSA can both be regarded as forms of normalisation layers and we can apply a re-weighting mechanism after them:
@@ 2.3 Attention Re-weighting Mechanism
where $\mathbf{O}$ denotes a matrix where every element is equal to one  [^2] To prevent training instability due to the absence of positive values in the initial rows, the elements within first three rows of $\mathbf{O}$ were set to zero.. The mean effective value of $\mathbf{A}\otimes\mathbf{N}$ is one, leading to an expectation of effective value of $\mathbf{A}\otimes\mathbf{N}-\mathbf{O}$ being zero. The notation $\operatorname{ReLU}^{p}$ indicates that the output of the ReLU function is raised to the power of $p$, with $p$ being a positive integer. ReLU effectively masks tokens with small attention scores, while the power operation enhances the discriminability of the remaining tokens. The re-weighting mechanism is applied post-training, requiring no additional gradient updates, which preserves the pre-trained model’s parameters intact and avoids the computational burden of fine-tuning.
As the token length increases, the attention scores become smoother and lacks distinct peaks, which hinders the model’s ability to establish connections between relevant tokens, thereby affecting the length extrapolation capabilities of transformers (Chiang & Cholak, 2022). This issue can be effectively addressed by the power operation, as demonstrated below:
@@
Let $x_{1},x_{2},\ldots,x_{n}$ represent the positive elements in one row of the ReLU output from Equation 5, with the maximum value denoted as $x_{m}$. The formula for any input $x_{j}$ in Equation 5 is:
$\overline{x}_{j}=\frac{x_{j}^{p}}{\sum_{k=1}^{n}x_{k}^{p}}$ (6)
For any $x_{l}$ where $x_{l}<x_{m}$, the distance between their re-weighted values with respected to $p_{\to}\infty$ is given by:
$\lim_{p_{\to}\infty}{\Delta(\overline{x}_{m}-\overline{x}_{l})}$ $=\lim_{p_{\to}\infty}{\frac{x_{m}^{p}-x_{l}^{p}}{\sum_{k=1}^{n}x_{k}^{p}}}$ (7)
$=\lim_{p_{\to}\infty}{\frac{1-(\frac{x_{l}}{x_{m}})^{p}}{\sum_{k=1}^{n}(\frac{x_{k}}{x_{m}})^{p}}}$
$=1$
@@
This limit demonstrates that as $p$ increases, the distance between the maximum value and any smaller value approaches 1. Specifically, when $p$ is large, the term $(\frac{x_{l}}{x_{m}})^{p}$ approaches zero for any $x_{l}<x_{m}$, causing all non-maximum values to be effectively suppressed. Meanwhile, the denominator becomes dominated by $x_{m}^{p}$, resulting in the maximum value approaching 1 while all other values approach 0. This property ensures that the attention mechanism maintains sharp, distinct peaks even with longer sequences, thereby preserving the model’s ability to focus on the most relevant tokens. We will refer to the combination of the re-weighting mechanism and LSSA as LSSAR in the rest of this paper.
3 Experiments
@@
The experiments were executed using 8 NVIDIA A100 80GB GPUs. We utilised the GPT-2 small architecture, consisting of 124 million parameters (Radford et al., 2019). Importantly, we substituted the original absolute position embeddings with Rotary Position Embeddings (RoPE) (Su et al., 2024) and incorporated NTK Scaling (Liu et al., 2024b) to handle extended token lengths during inference for length extrapolation.
@@
All models were trained using a sequence length of 1024 on the fineweb-10B dataset (Penedo et al., 2024), which consists of 10.2 billion training tokens distributed across 18,865 training steps, along with 0.1 billion validation tokens.
3.1 Softmax Decomposition
@@
As illustrated in Equation 2, the Softmax operation can be decomposed into a non-linear transformation via the exponential function, followed by the $l_{1}$-norm. In this section, we examine the importance of each component for LLMs by training the GPT-2 model from scratch with and without each component.
@@
The results are presented in Table 2. Comparing the model employing Softmax in Scenario IV with Scenario I, which does not utilise Softmax, reveals that performance does not significantly deteriorate without Softmax. This can be attributed to the masking operation in attention, which elevates the rank of raw attention scores from $d$ to $L$. Given that $L\gg d$, causal masking substantially enhances the model’s non-linear capability. In tasks devoid of causal masking, such as transformers used in vision applications, performance degrades markedly (Wortsman et al., 2023). Scenario II shows that using a non-linearity enhances performance. However, it is possible that alternative activation functions may enhance performance beyond that produced by using the exponential function.
@@
3.2 Comparison with Different Activation Functions
We modify the standard attention mechanism by introducing several novel attention variants for comparative analysis. Specifically, these variants are created by substituting $\phi$ in Equation 3 with several widely-used activation functions: ReLU (Rumelhart et al., 1986), ReLU2 (So et al., 2021), ReLU6 (Howard et al., 2017), GeLU (Hendrycks & Gimpel, 2016), Sigmoid (Rumelhart et al., 1986), Softplus (Zheng et al., 2015), and Mish (Misra, 2019).
@@ 3.2 Comparison with Different Activation Functions
The results presented in Table 3 indicate that all activation functions, except for Softplus, result in higher validation loss values compared to the standard attention mechanism represented by $e^{x}$. This phenomenon can be attributed to the distribution of positive values in the row attention scores (the inputs to these activation functions), which are predominantly concentrated around the diagonal, while other positions remain negative. Consequently, these activation functions, apart from Softplus, tend to suppress negative inputs, leading to a limited number of activated neurons during training. This suppression results in higher loss values than those observed with the exponential function, $e^{x}$. Nevertheless, the performance differences are relatively minor, suggesting that the positive values play a more critical role than the negative values in influencing the model’s output.
@@
Sigmoid (Ramapuram et al., 2024) 3.1935 7.4554 11.8355 14.4995
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}}$) 3.1930 6.5830 8.4679 9.5939
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.1849 4.3470 5.5544 6.1846
Sigmoid (Proposed) 3.2000 4.3811 5.8465 6.5701
ReLU (Wortsman et al., 2023) 3.2143 6.2662 8.4982 10.3460
ReLU (Li et al., 2022) 3.2006 4.5192 5.6924 6.4561
ReLU (Shen et al., 2023) 3.2155 6.5573 8.9072 10.7266
@@
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.1841 3.7387 4.9286 5.7450
@@
Sigmoid (Ramapuram et al., 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.2171 3.3499 3.6108 3.8587

### v4

@@ Abstract
[…]
Introduction
@@
Self-attention has been primarily responsible for the recent success of Large Language Models (LLMs). Self-attention enables models to assess the importance of different words within a sentence, capturing complex relationships and dependencies in the data. Its effectiveness is mainly attributed to the Softmax operation, as when the Softmax operation is omitted or replaced with alternative activation functions, model performance tends to decline (Wortsman et al. 2023; Ramapuram et al. 2024; Shen et al. 2023). Although widely used, Softmax self-attention has two main limitations. Firstly, it suffers from numerical instability due to the natural exponential function ($e^{x}$), especially when scaling model sizes to trillions of parameters (Qi et al. 2024). Secondly, as the token length increases during inference, the attention scores calculated by self-attention become smoother and lack distinct peaks. This ”attention smoothing” hinders the model’s ability to establish connections between relevant tokens, thereby crippling the length extrapolation capabilities of transformers (Chiang and Cholak 2022; Veličković et al. 2024). This paper addresses these two problems simultaneously by proposing a new design principle for attention mechanisms that offers improved numerical stability and dramatically better performance at large token lengths. Our core proposal is a new attention architecture designed as a two-stage process: a Normalisation Stage followed by a Sharpening Stage.
To develop a better normalisation stage, we first analysed the standard Softmax function. We deconstructed it into its two functional components: a non-linear positivity transformation ($e^{x}$) and a subsequent $l_{1}$-normalisation. Our experiments revealed that the $l_{1}$-normalisation is the critical component for maintaining model performance. This insight frees us to redesign the normalisation process for better stability and performance. We replace the exponential function with the more numerically stable Softplus activation and incorporate a dynamic length scale factor based on invariance entropy. This creates our new normalisation stage, a novel mechanism we call Length Scaled Softplus Attention (LSSA), which outperforms the standard attention mechanism not only at the training sequence length but also for longer sequences.
However, a normalisation stage alone does not solve the attention smoothing problem. Therefore, we introduce a second, explicit architectural stage: a re-weighting mechanism that sharpens the attention distribution. Applied after the LSSA normalisation, this stage amplifies critical token relationships and suppresses noise using a power transformation. This inherently sharpens attention peaks without needing post-hoc fixes like positional interpolation (Chen et al. 2023; Li et al. 2023), which retroactively stretches embeddings but fails to address the root cause of attention smoothing.
[…]
@@
  • We introduce a novel Softmax-free attention mechanism, LSSA, which incorporates a dynamic length scale factor based on invariance entropy, demonstrating superior performance compared to standard attention mechanisms across a variety of sequence lengths.
  • We propose a novel architectural re-weighting mechanism that fundamentally redesigns how attention scores are computed, inherently sharpening token relevance by amplifying critical weights and suppressing noise.
  • When LSSA is combined with the proposed re-weighting method, the resulting model (LSSAR) not only excels in length extrapolation while ensuring numerical stability—as demonstrated by its success in passkey retrieval tasks where standard attention fails—but also translates into superior performance in real-world applications.
Method
Operation Softmax Inverse Re-centred Re-centred & $l_{1}$-norm
Loss 3.1911 3.1915 3.2008 3.1954
Our proposed attention mechanism is designed as a two-stage phases. The first involves normalising the raw attention scores to create a stable distribution. The second re-weights this distribution to sharpen focus and enhance length extrapolation. Full details are provided in the following sub-sections.
Softmax Decomposition: The Normalisation Stage
@@
Scaled dot-product attention transforms queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) into an output. First, attention scores $\mathbf{A}$ are produced via Eq. 1. Here, $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{L\times d}$, where $L$ is the sequence length and $d$ is the dimensionality. The optional term $\mathbf{M}$ is a mask matrix of shape $L\times L$, which is essential in causal self-attention. The mask $\mathbf{M}$ is constructed with zeros on and below the diagonal and -$\infty$ above it, ensuring the attention mechanism only considers past and present tokens. The attention scores are then used to compute the output as $\mathbf{O}=\mathbf{A}\mathbf{V}$.
The Softmax function is a cornerstone of the attention mechanism. It is often assumed that its non-negative outputs are essential for good performance. However, this assumption is questionable, as replacing Softmax with other non-negative activation functions often leads to a decline in model performance (Wortsman et al. 2023; Ramapuram et al. 2024; Shen et al. 2023).
To investigate this, we conducted experiments to test the necessity of non-negative attention scores. As shown in Tab. 1, inverting the Softmax outputs by multiplying them with negative one resulted in a negligible change in validation loss, suggesting that non-negativity is not a critical property. We also re-centred the attention scores by subtracting the row-wise mean, creating a mix of positive and negative values [^1] To prevent training instability due to the absence of non-zero values in the initial rows, the first three rows were left unmodified.. This led to a minor increase in loss. However, when we subsequently applied $l_{1}$-normalisation to these re-centred scores, performance was almost fully restored.
These results indicate that the crucial component of Softmax is not its positivity but its normalisation property. Motivated by this, we decompose the Softmax operation into a non-linear positivity transformation, $\phi(\mathbf{x})=e^{\mathbf{x}}$, followed by an $l_{1}$-norm:
$\operatorname{Softmax}(\mathbf{x})$ $=\frac{e^{\mathbf{x}}}{\sum_{j}e^{x_{j}}}=\frac{e^{\mathbf{x}}}{\|e^{\mathbf{x}}\|_{1}}=\frac{\phi(\mathbf{x})}{\|\phi(\mathbf{x})\|_{1}}$ (2)
This allows us to generalise the attention mechanism’s normalisation stage as follows:
@@
The subscripts $i$ and $j$ represent the row and column indices respectively. $\otimes$ denotes element-wise multiplication. The mask $\mathbf{M^{\prime}}$ is constructed with ones on and below the diagonal and zeros above it. It is evident that the original attention function (Eq. 1) is a special case of this general form where $\phi(\mathbf{x})=e^{\mathbf{x}}$. This decomposition allows for greater flexibility in designing the normalisation stage by exploring alternative positivity functions $\phi$.
Length Scaled Softplus Attention
@@
Based on the general form of Eq. 3, we now introduce Length Scaled Softplus Attention (LSSA), a novel attention variant designed to enhance the scalability and performance of LLMs. LSSA introduces a novel scaling factor that accounts for both the sequence length and the model’s dimensionality, thereby addressing limitations associated with traditional attention methods in handling long sequences during inference.
@@
LSSA replaces the exponential function with the Softplus activation to introduce non-linearity while maintaining smooth gradients essential for stable training dynamics. Empirical testing of several widely-used activation functions in place of $\phi$ in Eq. 3, revealed that Softplus delivers the best performance (see Tab. 3), justifying its adoption.
@@
In contrast to scaled dot-product attention, cosine similarity attention employs the ll2-norm for each row of $\mathbf{Q}$ and $\mathbf{K}$. This approach has been shown to produce more moderate attention weights, which can enhance performance across various tasks (Henry et al. 2020; Dehghani et al. 2023; Liu et al. 2022). However, in high-dimensional spaces, as the number of dimensions increases, two randomly selected vectors are likely to become orthogonal. This phenomenon causes the elements of the product of normalised $\nicefrac{{\mathbf{Q}_{i}}}{{\|\mathbf{Q}_{i}\|_{2}}}$ and $\nicefrac{{\mathbf{K}_{i}^{T}}}{{\|\mathbf{K}_{i}\|_{2}^{T}}}$ to approach zero, thereby worsening the vanishing gradient problem. Consequently, a scale factor becomes essential for cosine similarity attention, and the factor should be associated with the dimensionality $d$.
@@
Previous work demonstrated that replacing the traditional scaling factor $\nicefrac{{1}}{{\sqrt{d}}}$ with $\nicefrac{{\log{L}}}{{\sqrt{d}}}$ in scaled dot-product attention enhances the length extrapolation capabilities of transformers (Chiang and Cholak 2022). Furthermore, (Su 2021) highlighted that the inclusion of the $\log{L}$ factor aids in maintaining entropy invariance with different token length, thereby facilitating better extrapolation to unknown sequence lengths. We extend this concept by introducing a dynamic length scale factor that adapts to the varying number of attended tokens in each row of the attention matrix. Specifically, we set the scaling factor to $\log{d}\log{\mathbf{N}}$, where $\mathbf{N}$ is an $L\times L$ matrix where each element in row $i$ is equal to $i$ (the number of tokens attended to in that row). This ensures the attention mechanism remains robust across varying sequence lengths. Hence, the formulation of LSSA is:
@@
Attention Re-weighting Mechanism: The Sharpening Stage
While the normalisation stage provides a stable attention distribution, it still activates all input tokens. This is counter-intuitive, as tokens early in a sequence are less likely to be relevant to the current token being processed. This leads to the attention smoothing problem (Chiang and Cholak 2022; Veličković et al. 2024), where attention scores lack distinct peaks as sequence length increases during inference.
Experiments with ReLU-based attention, shown in Tab. 3, demonstrate a minor performance decrease compared to the original attention mechanism. The observed performance decrement can be attributed to the distribution of positive values within the row attention scores. Specifically, positive values are predominantly concentrated around the diagonal, while other positions remain negative. The ReLU activation function tends to suppress these negative inputs, resulting in a limited number of activated neurons during training. This sparsity restricts the model’s capacity to capture essential relationships across the entire sequence, thereby adversely impacting overall performance.
This phenomenon cannot be adequately addressed by traditional normalisation methods such as LayerNorm (Ba, Kiros, and Hinton 2016), as LayerNorm does not alter the underlying distribution of row attention scores. One potential solution is to transform all attention score values to the positive domain using functions like $e^{x}$ or Softplus before applying LayerNorm. As all values become positive through such transformations, employing the $l_{1}$-norm emerges as a more effective strategy. The $l_{1}$-norm is advantageous because it is simple to calculate and ensures that the mean of the effective values in each row (the positive values that sum to one) remains constant (equals to $\nicefrac{{1}}{{i}}$), thereby maintaining the integrity of the attention distribution. This reframes both standard Softmax and our LSSA as a unified ’normalisation stage’. This conceptual separation allows us to introduce our re-weighting mechanism as a distinct second stage, applied after the initial normalisation:
@@
Here, the normalised attention scores $\mathbf{A}$ are first scaled by the token count $\mathbf{N}$ and then shifted by a matrix $\mathbf{O}$ (an offset matrix of ones, with zeros in the first three rows to prevent instability). This centres the distribution around zero. The $\operatorname{ReLU}^{p}$ function then masks scores below zero and sharpens the remaining positive scores by raising them to the power of a hyper-parameter $p$. A final $l_{1}$-normalisation ensures the output is a valid probability distribution. The re-weighting mechanism is an additional component of the attention layer, we refer to the combination of LSSA (stage 1) and this re-weighting mechanism (stage 2) as LSSAR in the rest of this paper.
@@
This power operation is key to solving the attention smoothing problem. To demonstrate this let $x_{1},x_{2},\ldots,x_{n}$ represent the positive elements in one row of the ReLU output from Eq. 5, with the maximum value denoted as $x_{m}$. The formula for any input $x_{j}$ in Eq. 5 is $\nicefrac{{x_{j}^{p}}}{{\sum_{k=1}^{n}x_{k}^{p}}}$. For any $x_{l}$ where $x_{l}<x_{m}$, the distance between their re-weighted values with respected to $p_{\to}\infty$ is given by:
$\overline{x}_{m}-\overline{x}_{l}=\lim_{p_{\to}\infty}{\frac{x_{m}^{p}-x_{l}^{p}}{\sum_{k=1}^{n}x_{k}^{p}}}=\lim_{p_{\to}\infty}{\frac{1-(\frac{x_{l}}{x_{m}})^{p}}{\sum_{k=1}^{n}(\frac{x_{k}}{x_{m}})^{p}}}=1$ (6)
@@
This limit demonstrates that as $p$ increases, the distance between the maximum value and any smaller value approaches 1. Specifically, when $p$ is large, the term $(\nicefrac{{x_{l}}}{{x_{m}}})^{p}$ approaches zero for any $x_{l}<x_{m}$, causing all non-maximum values to be effectively suppressed. Meanwhile, the denominator becomes dominated by $x_{m}^{p}$, resulting in the maximum value approaching 1 while all other values approach 0. This property ensures that the attention mechanism maintains sharp, distinct peaks even with longer sequences, thereby preserving the model’s ability to focus on the most relevant tokens.
Experiments
@@
The experiments were conducted using 8 NVIDIA A100 80GB GPUs. We utilised the GPT-2 small architecture (124 million parameters; Radford et al. 2019), replacing the original absolute position embeddings with Rotary Position Embeddings (RoPE; Su et al. 2024) and incorporating Dynamic NTK Scaling (emozilla 2023) to handle extended token lengths during inference.
@@
All models were trained using a sequence length of 1024 on the fineweb-10B dataset (Penedo et al. 2024), which consists of 10.2 billion training tokens distributed across 18,865 training steps, along with 0.1 billion validation tokens.
Softmax Decomposition
@@
As illustrated in Eq. 2, the Softmax operation can be decomposed into a non-linear transformation ($e^{x}$), followed by the $l_{1}$-norm. In this section, we examine the importance of each component for LLMs by training the GPT-2 model from scratch with and without each component.
@@
The results are presented in Tab. 2. Comparing the model employing Softmax in Scenario IV with Scenario I, which does not utilise Softmax, reveals that performance does not significantly deteriorate without Softmax. This can be attributed to the masking operation in attention, which elevates the rank of raw attention scores from $d$ to $L$. Given that $L\gg d$, causal masking substantially enhances the model’s non-linear capability. In tasks devoid of causal masking, such as transformers used in vision applications, performance degrades markedly (Wortsman et al. 2023). Scenario II shows that using a non-linearity enhances performance. However, it is possible that alternative activation functions may enhance performance beyond that produced by using the exponential function.
@@
Comparison with Different Activation Functions
We modify the standard attention mechanism by introducing several novel attention variants for comparative analysis. Specifically, these variants are created by substituting $\phi$ in Eq. 3 with several widely-used activation functions: ReLU (Rumelhart, Hinton, and Williams 1986), ReLU2 (So et al. 2021), ReLU6 (Howard et al. 2017), GeLU (Hendrycks and Gimpel 2016), Sigmoid (Rumelhart, Hinton, and Williams 1986), Softplus (Zheng et al. 2015), and Mish (Misra 2019).
@@
Sigmoid (Ramapuram et al. 2024) 3.1935 7.4554 11.8355 14.4995
Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}}$) 3.1930 6.5830 8.4679 9.5939
Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.1849 4.3470 5.5544 6.1846
Sigmoid (Tab. 3) 3.2000 4.3811 5.8465 6.5701
ReLU (Wortsman et al. 2023) 3.2143 6.2662 8.4982 10.3460
ReLU (Li et al. 2022) 3.2006 4.5192 5.6924 6.4561
ReLU (Shen et al. 2023) 3.2155 6.5573 8.9072 10.7266
@@
Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.1841 3.7387 4.9286 5.7450
@@
Sigmoid (Ramapuram et al. 2024)($b=-\log{\mathbf{N}},l_{1}$-norm) 3.2171 3.3499 3.6108 3.8587
@@
The results presented in Tab. 3 indicate that all activation functions, except for Softplus, result in poorer validation loss values compared to the standard attention mechanism represented by $e^{x}$. This phenomenon can be attributed to the distribution of positive values in the row attention scores (the inputs to these activation functions), which are predominantly concentrated around the diagonal, while other positions remain negative. Consequently, these activation functions, apart from Softplus, tend to suppress negative inputs, leading to a limited number of non-zero attention scores during training. This suppression results in higher loss values than those observed with $e^{x}$. Nevertheless, the performance differences are relatively minor, suggesting that the positive values play a more critical role in influencing the model’s output.